Jarvene Corporation uses the FIFO method in its process costing system. The following data are for the most recent month of operations in one of the company’s processing departments:

Units in beginning inventory 400
Units started into production 4,330
Units in ending inventory 280
Units transferred to the next department 4,450
Materials Conversion
Percentage completion of beginning inventory 80 % 20 %
Percentage completion of ending inventory 80 % 30 %

The cost of beginning inventory according to the company’s costing system was $7,847 of which $4,879 was for materials and the remainder was for conversion cost. The costs added during the month amounted to $180,814. The costs per equivalent unit for the month were:

Materials Conversion
Cost per equivalent unit $18.00 $23.00

Required:

1. Compute the total cost per equivalent unit for the month.

2. Compute the equivalent units of material and conversion in the ending inventory.

3. Compute the equivalent units of material and conversion that were required to complete the beginning inventory.

4. Compute the number of units started and completed during the month.

5. Compute the cost of ending work in process inventory for materials, conversion, and in total for the month.

6. Compute the cost of the units transferred to the next department for materials, conversion, and in total for the month.
Answer #1

1) Total Cost per equivalent unit = 18+23 = $41 per unit

2) Equivalent unit in ending inventory

Material Conversion
Equivalent unit of ending inventory 280*80% = 224 280*30% = 84

3) Equivalent unit of beginning inventory

Material Conversion
Equivalent unit of beginning inventory required to complete 400*20% = 80 400*80% = 320

4) Unit started and completed = 4450-400 = 4050 Units

5) Calculate cost of ending work in process

Material Conversion Total
Total Cost of ending work in process 224*18 = 4032 84*23 = 1932 5964

6) Calculate cost of unit transferred out

Material Conversion Total
Cost of units completed and transferred out (4130*18+4879) = 79219 (4370*23+2968) = 103478 182697